• Each month, ​
  • 1% of employed workers lose their jobs (s = 0.01)​
  • 19% of unemployed workers find jobs (f = 0.19)​
  • Find the natural rate of unemployment:
1
Expert's answer
2021-01-22T09:41:43-0500

"\frac{U}{L}=\frac{s}{s+f}=\frac{0.01}{0.01+0.19}=0.05, or" "5\%"